OpenStudy (anonymous):

I've had some difficulty with equation of the tangent line. F(X) = (-10x^2-3)/(4x+1) ; x = 0

OpenStudy (anonymous):

Are you trying to find the equation of the tangent line?

OpenStudy (anonymous):

yes i am. equation of the tangent line to the graph y = f(x) at the point on the graph where x has the inidcated valued

OpenStudy (anonymous):

Ok. So you must find the slope of this line first, correct?

OpenStudy (anonymous):

indeed, do i just plug it into the equation to find x then, find the derative, then plug it into the slope formula? y-y1 = m(x-x1)

OpenStudy (anonymous):

well you must find the derivative of the original equation to find the slope of the tangent line. after derivating the original equation, plug in x=0. that will give you the slope of your tangent line at x=0

OpenStudy (anonymous):

would f(x) be -3/1 at the slope when x =0

OpenStudy (anonymous):

lemme do it..gimme a sec plz

OpenStudy (anonymous):

Hmm, I got 12

OpenStudy (anonymous):

let me do the problem again, generally i just did a quick overview of one part, but thank you! i appreciate it.

OpenStudy (anonymous):

I'll do it again also

OpenStudy (anonymous):